Consider the circuit below. R1 = 3 ohms, R2 = 6 ohms, R3 = 2 ohms, R4 = 4 ohms, and the battery has a voltage of 12 V. Answer the following

R3 R2- R1. R4.

Pick the best answer for each True/False answer

1. R1 and R2 will both have a voltage across each of 12V. True or False.

2. R1 and R2 will both have the same current traveling through them. True or False.

3. R3 and R4 will both have a voltage across each of 12 V. True or False.

4. R3 and R4 will both have the same current traveling through them. True or False.

5. Calculate the current through R1.

6. Calculate the voltage across R3 and R4.

7. Calculate the total resistance of the circuit.
Answer #1

Re: 3 h. R2= 62, Rz: 200, Ry: 44, V: 12v. (1) (2). (3) 14). (5) True, sen parallal voltage is same) False. I current will not
Answer #2

image.png

image.png

image.png

source: My Physics Knowledge
answered by: Sadaf Mukhtar
Answer #3

ANSWER :


1. True


2. False


3. False.


4. True


5. 


I1 = Current through R1 = V/R1 12/3 = 4 amps ( ANSWER)


6.


I3 = Current through R3 and R4 = V/((r3 + R4) = 12/(2 + 4) = 2 amps.


So,


V3 = Voltage across R3 = I3 R3 = 2 * 2 = 4 Volts (ANSWER).


V4 = Voltage across R4 = I3 R4 = 2 * 4 = 8 Volts (ANSWER).


6.


I2 = current through R2 = 12/6 = 2 amps

I = Total current from the battery = I1 + I2 + I3 = 4 + 2 + 2 8 amps

Total resistance of the system, R = V/I = 12/8 = 1.5 Ω (ANSWER).

answered by: Tulsiram Garg